OpenStudy (anonymous):

Get a medal :3 Solve for x x-3/2 = -5 A. –13 B. –7 C. –16 D. –4

OpenStudy (anonymous):

\[\frac{ x -3 }{ 2}\]\[= -5\]

OpenStudy (anonymous):

I think it's -7 :3

OpenStudy (anonymous):

and you are completely correct, very nice job

OpenStudy (anonymous):

@Embryo Ok, thanks ^.^

OpenStudy (anonymous):

X-3/2=-5 (2)x-3=-5(2) X-3 = -10 X=-10+3 x=-7

OpenStudy (anonymous):

I wanted to show you, so you will know how to do it next time. :)